OpenStudy (anonymous):

verify cot(x)-tan(x)=4 cos^2(x)-2/sin(2x)

OpenStudy (anonymous):

Have you tried using trig identities and substitution?

OpenStudy (anonymous):

http://www.sosmath.com/trig/Trig5/trig5/trig5.html Here are some good ones to refer to. Remember that 'u' can be the same as 'x'. It's just a variable. :)

OpenStudy (anonymous):

yeah i tried using them but i can't get throught problem so that the left side equals the right side, I NEED HELP!

OpenStudy (anonymous):

this is what i first did: i turn\[\cot (x) - \tan (x) \] to \[\frac{ \cos(x) }{ \sin(x) } - \frac{ \sin(x) }{ \cos(x) }\]

OpenStudy (anonymous):

and then i multiplied the left side by \[\frac{ \cos(x) }{ \cos(x) }\]

OpenStudy (anonymous):

and the right hand side by\[\frac{ \sin(x) }{ \sin(x) }\]

OpenStudy (anonymous):

to get: \[\frac{ \cos^2(x) }{ \cos(x)\sin(x) } - \frac{ \sin^2(x) }{ \cos^2(x)\sin^2(x) }\] and then that simplifies to : \[\frac{ \cos^2(x)-\sin^2(x) }{ \cos(x)\sin(x) }\]

OpenStudy (anonymous):

any further help??

terenzreignz (terenzreignz):

Do you still need any?

OpenStudy (anonymous):

Seems like @terenzreignz has solved your query mostly. The numerator becomes cos2x and denominator becomes sin(2x)/2

terenzreignz (terenzreignz):

I haven't done anything o.O Give credit where credit is due, @NeetziD

OpenStudy (anonymous):

Telling em that they already have what they're looking for... that needs an eye. :\

terenzreignz (terenzreignz):

Gotcha.

OpenStudy (anonymous):

wait what?...

terenzreignz (terenzreignz):

I was asking if you needed any more help with this.